SPECTROSCOPY AND THE DISCOVERY OF THE FULLERENES

\underline{R. F. CURL}, Department of Chemistry and Rice Quantum Institute, Rice University, Houston, TX 77005.

Spectroscopy played a small but interesting role in motivating the work that led to the discovery of the fullerenes and in our subsequent work on these compounds. These aspects will be fitted into the story of the fullerene discovery and its aftermath.